OpenStudy (anonymous):

Three condesers of 6,10 and 15uf are connected in series across 200v supply.If the condensers are disconnected,how much energy can the condensers charged if they are connected in parallel?

OpenStudy (anonymous):

im not sure , its my calculation Energy = 1/2 CV^2 if parallel then Ce = C1 + C2 + C3

OpenStudy (anonymous):

is it 31uf is 31x10^-6 ? so (31x10^-6)(200v)/2 ?

OpenStudy (anonymous):

E =( 31x10^-6)(200)^2 / 2

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

I think thats it but the first part if in series confusing me . Need a second opinion

OpenStudy (radar):

Total capacitance for series connected capacitors is computed in a similar manner as resistors connected in parallel. In other words the total capacitance is only 3 microfarad for the above arrangement.